OpenStudy (anonymous):

A right triangle has a hypotenuse of length 21 and a leg of length 14. What is the length of the other leg of the triangle?

OpenStudy (nottim):

why not use pythagorean theorem?

OpenStudy (nottim):

c^2=a^2+b^2

OpenStudy (nottim):

but rearrange to solve for a

OpenStudy (nottim):

that's all!

OpenStudy (anonymous):

Do I have to find the length of the other side first?

OpenStudy (anonymous):

@NotTim

OpenStudy (anonymous):

21^2-14^2 = 245 Square root of 245 = 7(sqrt)5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You're correct.

OpenStudy (anonymous):

Thank you!